OpenStudy (anonymous):

Find the GCF of the pair of monomials. 14xy^2 and 21y^3

OpenStudy (anonymous):

What is the biggest coefficient that will go into both 14 and 21?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

Ok, now how many y's can go into both y^2 and y^3?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

To be accurate, y^2 goes into y^2 once, and y^3 "y times", because y^2 * 1 = y^2 and y^2 * y = y^3. So your GCF is 7y^2. In factored form this would look like: \[7y^2 (2x + 3y)\]

OpenStudy (anonymous):

but what about the x

OpenStudy (anonymous):

It's in there. You cannot factor the x because both terms don't have an x.

OpenStudy (anonymous):

o ok thank you

OpenStudy (anonymous):

No problem.